OpenStudy (anonymous):

<1 and <2 are supplementary angles. m <1 is 4y + 7 and m <2 is 9y + 4. Find m.

OpenStudy (anonymous):

4y + 7 + 9y + 4 = 180 degrees

OpenStudy (anonymous):

combine like terms and solve for y. 13y + 11 = 180 13y = 169 y = 13 Angle 1 = 4(13) + 7= 59 Angle 2 = 9(13) + 4 = 121